Part 1


An elevator has a placard stating that the maximum capacity is 1376 lb—8 passengers.​ So, 8 adult male passengers can have a mean weight of up to 1376/8=172 pounds. If the elevator is loaded with 8 adult male​ passengers, find the probability that it is overloaded because they have a mean weight greater than 172 lb.​ (Assume that weights of males are normally distributed with a mean of 178 lb and a standard deviation of 27 lb​.) Does this elevator appear to be​ safe?


Question content area bottom


Part 1


The probability the elevator is overloaded is



.


1
Expert's answer
2021-12-29T17:16:17-0500

Given : Mean "(\mu)=174" , standard deviation "(\sigma)=28" , sample size "(n)=8"

Formula: "z=\frac{\bar{x}-\mu}{\frac{\sigma}{\sqrt{n}}}"

Find: "P\text{( sample mean greater than 172)}=P(\bar{x}>172)=?"

"\begin{aligned}

&\Rightarrow P\left(\frac{\bar{x}-\mu}{\frac{\sigma}{\sqrt{n}}}>\frac{172-174}{\frac{28}{\sqrt{8}}}\right) \\

&\Rightarrow P(z>-0.2020)

\end{aligned}"

Note: "P(z>a)=1-P(z<a)"

"\Rightarrow 1-P(z<-0.2020)"

Refer Standard normal table/Z-table to find the probability or use excel formula "=NORM.S.DIST(-0.2020, TRUE)" to find the probability.

"\begin{aligned}

&\Rightarrow 1-0.4200 \\

&\Rightarrow {0 . 5 8 0 0}

\end{aligned}"

The probability the elevator is overloaded is 0.5800

No, there is a good chance that 8 randomly selected people will exceed the elevator capacity.